OpenStudy (anonymous):

Solve equation using the quadratic formula. x^2+25=0. Please show me the steps.

OpenStudy (anonymous):

is this dealing with imaginary numbers?

OpenStudy (anonymous):

we havn't done with imaginary numbers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x^2+25 (x+5)(x-5)=0 x=-5, x=5

OpenStudy (anonymous):

Is it difference between the squares

OpenStudy (anonymous):

Do u think we need not to use the quadratic formula

OpenStudy (anonymous):

Add 25 to both sides of the equation to get x^2 = 25 Take the square root of both sides of the equation to get x = ±√25 = ±5 x = 5 and x = -5 are solutions to this equation. You can recognize that the difference of two squares is a standard form that you know the factors for. A2 - b2 = (a+b)(a-b) x2 - 25 = (x+5)(x-5) = 0 Values of x that make the factors zero are x = -5, x = 5 (same as above)

OpenStudy (anonymous):

Does this help?

OpenStudy (anonymous):

Yes .Thank you so much.

OpenStudy (anonymous):

Anytime :-)